A 10-Newton object moves to the left at 1 m/s. Its kinetic energy is approximately _____ Joules.

 

a. 0.5

b. 1

c. 10

d. more than 10


1
Expert's answer
2022-02-28T10:42:59-0500

The kinetic energy

"E_k=\frac{mv^2}{2}=\rm\frac{1\: kg*(1\: m/s)^2}{2}=0.5\:J"

Answer: a